# The DeepSpeak Dataset

### v3

@@ Appendix A Prompts
A.1 Video Action Prompts
@@ 5. If I could time travel, where in the past or future would I like to go?
Table 3: Video action prompts.
A.2 Standardized Scripted Prompts
We apply auditory modeling to computer speech recognition.
He sank back sighing and was soon asleep again.
The mango and the papaya are in a bowl.
Will you tell me why.
That is what childhood is he told himself.
As a rule part time farmers hire little help.
The new birth is miraculous and mysterious.
The fear of punishment just didn’t bother him.
The figure in the corner belched loudly a deep liquid eruption.
Table 4: Standardized scripted prompts, consisting of ten separate sentences for version 1.0 and 2.0.
A.3 Voice Cloning Prompts
The birch canoe slid on the smooth planks.
Glue the sheet to the dark blue background.
It’s easy to tell the depth of a well.
These days a chicken leg is a rare dish.
Rice is often served in round bowls.
The juice of lemons makes fine punch.
The box was thrown beside the parked truck.
The hogs were fed chopped corn and garbage.
Four hours of steady work faced us.
A large size in stockings is hard to sell.
Table 5: Voice cloning input prompts for version 1.0 and 2.0, consisting of ten consecutive sentences and one continuous paragraph.
A.4 Randomized Scripted Prompts

### v4

@@ Appendix I Visual Identity Pairing Examples
Q.4 Video Action Prompts
Table 12: Video action prompts.
@@ 5. If I could time travel, where in the past or future would I like to go?
Q.5 Example frames from action prompts
read wave hand turn head right/up tilt head down lean in
Figure 26: Representative examples of different action prompt types.
Q.6 Randomized Scripted Prompts
@@
Appendix R Video Identity Matching
Table 14: Video identity pairs found through the minimal cosine distance between mean CLIP embeddings of five video frames. Part 1/2.
identity_1 identity_2
0 1
2 3
4 5
6 7
8 9
10 11
12 13
14 15
16 17
18 19
20 21
22 23
24 25
26 27
28 29
30 31
32 33
34 35
37 38
39 40
41 42
43 44
45 46
47 48
49 50
51 52
53 54
55 56
57 58
59 60
61 62
63 64
65 66
67 68
69 70
71 72
73 74
75 76
77 78
79 80
81 82
83 84
identity_1 identity_2
85 86
87 192
88 126
89 106
90 194
91 119
92 184
93 207
94 191
95 201
96 142
97 98
99 100
101 174
102 196
103 134
104 193
105 177
107 170
108 109
110 111
112 113
114 115
116 124
117 208
118 165
120 129
121 214
122 137
123 162
125 220
127 172
128 164
130 152
131 149
132 206
133 173
135 205
136 219
138 145
139 185
140 141
identity_1 identity_2
143 144
146 171
147 154
148 155
150 151
153 163
156 157
158 159
160 161
166 167
168 178
169 202
175 198
176 197
179 218
180 181
182 183
186 215
187 217
188 211
189 190
195 221
199 200
203 204
209 210
212 213
222 258
223 419
224 230
225 411
226 270
227 400
228 480
229 253
231 414
232 361
233 462
234 264
235 393
236 467
237 383
Table 15: Video identity pairs found through the minimal cosine distance between mean CLIP embeddings of five video frames. Part 2/2.
identity_1 identity_2
238 408
239 243
240 323
241 428
242 256
244 461
245 301
246 247
248 274
249 443
250 407
251 271
252 474
254 445
255 315
257 310
259 332
260 473
261 334
262 280
263 307
265 442
266 449
267 272
268 450
269 298
273 459
275 313
276 370
277 398
278 309
279 423
281 427
282 367
283 312
284 377
285 339
286 384
287 388
288 422
289 487
290 415
identity_1 identity_2
291 305
292 417
293 485
294 478
295 421
296 369
297 340
299 346
300 379
302 416
303 446
304 344
306 397
308 316
311 390
314 406
317 434
318 319
320 402
321 350
322 345
324 409
325 412
326 396
327 457
328 381
329 440
330 454
331 436
333 418
335 490
336 426
337 472
338 405
341 373
342 372
343 356
347 452
348 441
349 471
351 401
352 463
identity_1 identity_2
353 382
354 359
355 386
357 376
358 439
360 447
362 420
363 453
364 438
365 375
366 378
368 425
371 404
374 380
385 482
387 429
389 392
391 479
394 448
395 475
399 444
403 481
410 458
413 431
424 486
430 476
432 455
433 477
435 469
437 470
451 456
460 483
464 465
466 489
468 484
488 491
492 496
493 494
495 497
498 500
499 501